OpenStudy (anonymous):

Help Medal The Great Pyramid of Giza is a square pyramid located by El Giza, Egypt. It is believed that the original height of the pyramid was 480.6 feet. The base is a square with side length of 755.9 feet. Find the volume of the original Great Pyramid of Giza. Explain your reasoning.

OpenStudy (anonymous):

I keep geting \[5.83\times10^{7}\]

OpenStudy (anonymous):

which multiplies into a crazy number that I dot think is realistic. Am I missing something or not doing it right please help.

Directrix (directrix):

Vpyramid = (1/3)* B * h where B is the area of the square base and h is the height. So, crank this out and see what you get: V = (1/3)* B * h V = (1/3) * (755.9) ^2 * (480.6) = ? @BrAnDoN_AAS